OpenStudy (anonymous):

What is the approximate distance from (6, 5) to (2, 1)?

OpenStudy (anonymous):

Hi Welcome to Openstudy! To calculate the distance between two points, you have to use the distance formula: \(\LARGE d=\sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2}\)

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

just substitute your points: \((x_1,y_1)=(6,5)\\(x_2,y_2)=(2,1)\)

OpenStudy (anonymous):

ohh ok thank you

OpenStudy (anonymous):

so it will be like this: \(d=\sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2}\\=\sqrt{(6 - 2)^2 + (5 - 1)^2}\\d=?\) you're welcome :)